# Assemblage: Automatic Binary Dataset Construction for Machine Learning

### v1

@@ Abstract
Binary code is pervasive, and binary analysis is a key task in reverse engineering, malware classification, and vulnerability discovery. Unfortunately, while there exist large corpuses of malicious binaries, obtaining high-quality corpuses of benign binaries for modern systems has proven challenging (e.g., due to licensing issues). Consequently, machine learning based pipelines for binary analysis utilize either costly commercial corpuses (e.g., VirusTotal) or open-source binaries (e.g., coreutils) available in limited quantities. To address these issues, we present Assemblage: an extensible cloud-based distributed system that crawls, configures, and builds Windows PE binaries to obtain high-quality binary corpuses suitable for training state-of-the-art models in binary analysis. We have run Assemblage on AWS over the past year, producing 890k Windows PE and 428k Linux ELF binaries across 29 configurations. Assemblage is designed to be both reproducible and extensible, enabling users to publish “recipes” for their datasets, and facilitating the extraction of a wide array of features. We evaluated Assemblage by using its data to train modern learning-based pipelines for compiler provenance and binary function similarity. Our results illustrate the practical need for robust corpuses of high-quality Windows PE binaries in training modern learning-based binary analyses. Assemblage can be downloaded from https://assemblage-dataset.net/.
@@ 1 Introduction
Binary analysis, the task of understanding and analyzing an executable binary program, is extraordinarily labor-intensive. Practitioners often spend hours to weeks manually analyzing a single program (Raff et al., 2020). The development of automated binary analysis tools has also proven to be extremely labor-intensive. Historically, such tools have relied on careful and judicious software engineering efforts to handle all corner cases, or used SAT solvers and similar data structures to resolve as many conditions as possible (Liu et al., 2023b). Given this high human effort, it is unsurprising that an increasing number of binary analysis methods use machine learning to help automate such efforts. This includes low-level tasks such as function disassembly (Yu et al., 2022), and high-level tasks like malware detection (Rokon et al., 2020) or variable name identification (Lacomis et al., 2019; Chen et al., 2022) which leverage these lower-level results.
@@
However, significant issues in data availability have hampered the development of machine learning-based binary analysis for decades (Abt and Baier, 2014; Raff and Nicholas, 2020). Disassembly and decompilation require a mapping between the original true source or assembly code and the corresponding bytes of binary instructions, but it is non-trivial to collect and maintain such a mapping. Malware detection requires access to a collection of “benign” programs to train on, but can not be legally distributed due to copyright and license concerns. The situation is particularly bad for Windows PE binaries: while Windows is the most frequent target of malware, modern binary analysis systems train almost exclusively on Linux ELF binaries. Toolchain identification requires knowledge of the entire compilation process and of a variety of compiler settings that are simply not recorded in the original file. Binary function similarity requires having ground truth of functions that are semantically equivalent but written or compiled differently; most existing models rely on coding competition data that does not reflect the diversity of real-world code. The ad-hoc approaches taken by prior works to tackle these issues further complicate the reproducibility of these methods due to the infrastructure challenges in setting up an environment to build the same tools and perform complex data pre-processing (Marcelli et al., 2022; Raff, 2019, 2021; Gundersen et al., 2022). Moreover, authors often do not have the necessary licenses or do not make the code publicly available for others to extend the work.
@@
We introduce Assemblage, a tool that makes significant progress towards overcoming all of these challenges. Assemblage is a framework for crawling code hosting platforms, downloading open-source projects, compiling them with multiple different compiler versions and settings, and recording the necessary information to reconstruct byte-level mappings from the output back to the source code. By using open-source code targets from GitHub, we obtain greater diversity in programs, resolve licensing/distribution issues, and provide detailed ground truth to support all of the aforementioned research areas. This alleviates the concerns with non-semantic preserving modifications (Fleshman et al., 2018) and the exorbitantly expensive cost and more limited diversity of binary re-writing (Wong et al., 2023; Boutsikas et al., 2021). Assemblage is designed to be extensible so that additional sources of code, compilers, and feature extractors can be added for further functionality. Assemblage datasets can be further distributed as a list of repositories and settings so that issues with mixed-license repositories (e.g., GPL and BSD) do not make a binary-based distribution unlicensable. Finally, we have designed and tested the system to ensure the reproducibility of datasets built by Assemblage.
Modern binary analysis systems use ever-complex models, with the recent switch to large-language models (LLMs) necessitating orders-of-magnitude more data given the rapidly increase in network size and number of parameters. In Section 2 we survey several of these efforts, observing that the largest system trains only 48k Linux ELF binaries. We designed Assemblage to produce large, diverse corpuses of binaries for training these state-of-the-art models at unprecedented scale. Our experiments running Assemblage over the past year on AWS has produced a large corpus of benign Windows PE (890k) and Linux ELF (428k) binaries. Assemblage clearly tracks source-to-binary provenance (down to the instruction level, in the case of Windows PE binaries) to provide a rich array of features and facilitate compliance with licensing and distribution. Assemblage’s design facilities a wide degree of customization and extension, allowing almost every aspect of its pipeline to be customized to suit the needs of modern, data-intensive binary analysis tools.
@@
  • Assemblage, a cloud-based distributed system carefully engineered to automatically construct diverse corpuses of high-quality Windows PE binaries while ensuring compliance with licensing and terms of service.
  • An evaluation of Assemblage’s scalability and reproducibility, generating over 890k binaries over 20 weeks.
  • The Assemblage Windows and Linux datasets, consisting of 890k and 428k binaries respectively, with a total of 29 different combinations of CPU architecture, optimization level, compiler, compiler version, and other compiler flags.
  • Three case studies that demonstrate Assemblage’s application to (a) compiler provenance detection and (b) function similarity identification. Our case studies demonstrate the need for Assemblage’s rich features, particularly for corpuses of Windows PE binaries.
2 Related Work and Open Challenges
Prior Work: Crawling Open Source Repositories
Table 1: Binary corpuses—Assemblage vs. the literature
Dataset Binaries Functions Toolchain Dataset Format Public Assemblage-WinPE 890k 252M MSVC Binaries + PDB Assemblage-WinPE-Licensed 97k 41M MSVC Binaries + PDB Assemblage-LinuxELF 428k 316M GCC/Clang Binaries Assemblage-LinuxELF-Licensed 211k 186M GCC/Clang Binaries EMBER 1.1M undisclosed undisclosed Extracted features Features PassTell 552 149k GCC/Clang/ICC Extracted features DIRE 164k 3M GCC Binaries SPEC(XDA) 3.1k 9.75k GCC/ICC/MSVC Custom format GHCC (DIRTY) 75k  1M GCC Binaries VirusShare 67M undisclosed undisclosed Binaries BinaryCorp-26M(jTrans) 48k 25.8M GCC Binaries SOREL-20M 10M undisclosed undisclosed Binaries
Significant prior work has focused on crawling open-source repositories to build artifacts. For example, Sourcer (Linstead et al., 2007) crawled SourceForge to create a corpus of software for source-code similarity. More recent work targets GitHub due to its current dominance. This includes projects such as GHTorrent, which crawls and indexes GitHub (Gousios, 2013), along with other work focusing on improving crawling technology  (Barmpis et al., 2020). The open-source tool GHCC builds Linux-based GitHub repositories in batch mode and supports several other systems (such as DIRTY, XDA, and DIRE) (Lacomis and Hu, 2020).
Assemblage distinguishes itself from prior work by offering multi-node scalability, end-to-end reliability, an extensible interface, and deep integration with Windows toolchains, and a comparison is shown in Table 1. To the best of our knowledge, there is no preexisting work that provides a systematic and replicable solution to generate Windows binaries from a wide range of sources. In fact, much of the recent work applying machine learning techniques to binary analysis tasks (e.g. (Du et al., 2022; Wang et al., 2022)) relies on custom crawlers that cannot be extended to Windows PE binaries due to GUI-based toolchains.
@@ 2 Related Work and Open Challenges
The challenges of obtaining and distributing benign Windows PE executables continues to push modern work to rely upon manually-curated corpora of Linux binaries, or, alternatively, on paid services. For those who can afford it, VirusTotal has become the de-facto standard in malware classification due to its offering of benign binaries (Salem et al., 2021; Zhu et al., 2020b, a). For those who cannot, there are very few other options. While the distribution of malicious binaries is unfettered by licensing restrictions (e.g., VirusShare by Roberts (2011)), these restrictions can complicate distribution of benign corpuses of PE binaries, subsequently inhibiting reproducibility of machine-learning-based binary analysis tools. Additionally, even when data is available, it is often in preprocessed form (e.g., weights or features) that precludes its application to novel tasks. For example, while the popular EMBER data set includes 300k malicious and 300k benign Windows PE binaries, only EMBER’s features are public, not the binaries themselves (Anderson and Roth, 2018). While much work uses open-source Linux binaries (e.g., DIRE by Lacomis et al. (2019)), our case studies show that the utility of Linux-based training data for Windows-based tasks is unclear.
Consistent with our description of the landscape, there were many systems to which we were unable to apply Assemblage, due primarily to the lack of (a) available comparison data or code, or (b) support for Windows PE binaries. For example, work such as XDA makes data available only in preprocessed form (Pei et al., 2021). Similarly, tools such as PalmTree (a model for learning instruction embeddings) do not include implementations of the downstream tasks necessary to perform end-to-end comparison (Li et al., 2021b). Finally, almost all systems (such as Debin (He et al., 2018), Symlm (Jin et al., 2022), and DIRTY (Chen and Sun, 2022)) work only on Linux binaries, relying upon DWARF or ELF-specific details for feature extraction. We speculate that the development of machine-learning-based binary analysis systems is significantly hampered by the serious costs involved with obtaining large benign corpora of PE binaries.
Models Continue to Grow; Binary Datasets Must Keep Up
The most recent developments in binary analysis have been driven by the success of Large Language Models (LLMs) for natural language tasks. Encouraged by the impressive performance of models like BERT (Devlin et al., 2019a) and GPT-4 (OpenAI, 2023), researchers have adapted the methods of natural language processing to apply to the domain of binary executables. One branch of work has focused on building general assembly language models, which can be fine-tuned to perform a wide variety of tasks (Li et al., 2021b; Artuso et al., 2022; Koo et al., 2021; Zhu et al., 2023). Other work presents models with specific downstream tasks in mind, including binary code similarity detection (Wang et al., 2022; Ahn et al., 2022; Gu et al., 2023; Liu et al., 2023a), compiler provenance identification (He et al., 2022), vulnerability detection (Yu et al., 2021), and malware detection (Li et al., 2021a; Hamad et al., 2021; Wu et al., 2021).
While these LLM-inspired models have advanced the state of the art in many areas of binary analysis, they demand orders-of-magnitude more input data to avoid overfitting, magnifying the effect of PE binary availability crisis. LLMs derive their name partially from the fact that they are trained on massive amounts of data, and it is well known that LLMs can achieve significantly better performance and even unlock emergent abilities at scale (Brown et al., 2020; Wei et al., 2022). However, a surprising amount of existing work applies LLMs to relatively small input corpuses. We were able to find dataset breakdowns for 10 of 13 models mentioned in the previous paragraph; the average number of binaries used to compose the training set for these models is roughly over 13k (median 3.7k). Two models are trained on fewer than 200 binaries, with Liu et al. (2023a) being trained on only 48 binaries. Wang et al. (2022) and Zhu et al. (2023) are trained with the BinaryCorp dataset, which, at 48,130 (Linux) binaries, is more than twice the size of the next largest dataset used. We hope that the orders-of-magnitude greater amount of data made available by Assemblage (890k Windows PE and 428k Linux ELF binaries) will further improve the performance of these state-of-the-art tools.
@@
We found that nearly half of the transformer-based models discussed above are trained largely, or even primarily, on executables from standard Linux packages such as coreutils, binutils, diffutils, findutils, busybox, libcurl and openssl. While these packages are well-written and easy to access, compile, and parse, they do not align with the quality and characteristics typically found in the codes written by the majority of programmers. The over-reliance on common Linux packages for training exemplifies a major problem with these training datasets; out of the thirteen transformer-based models discussed, only one, Wu et al. (2021), is trained using Windows data. Moreover, the reliance on small collections of Linux binaries extends beyond transformer-inspired models (Marcelli et al., 2022; Kim et al., 2022; Al-Kaswan et al., 2023).
@@
With the increasing use of binary analysis in security-relevant settings, the absence of publicly available corpora of benign Windows binaries is a critical concern. The difficulty of obtaining Windows PE data in academic literature is prevalent in related domains like malware detection that rely on small datasets (Joyce et al., 2022), effect sizes (Patel et al., 2023), and label noise (Joyce et al., 2021a, b, 2023), which generally depend on VirusShare and VirusTotal for data. An examination of files uploaded to VirusShare in 2023 indicates that 40.38% of the roughly 1.8 million malware samples are Windows binaries, whereas less than one percent are ELF binaries (Roberts, 2011). [^1] The other major file types are PDF, which accounts for roughly 25% of files, and Unknown, which accounts for 32% of files and captures files whose types could not be parsed. Files in the unknown category include data files, some scripts, ASC files, and source code. Given the prevalence of Windows executables, and the fact that Windows accounts for roughly seventy percent of global operating system market share (Statcounter, 2023), it is troubling that nearly all the recent literature on machine learning approaches to binary analysis uses Linux binaries as training data. This is especially worrying because, as we demonstrate in Sections 6.2 and 6.3, models trained on Linux binaries do not necessarily generalize to the Windows domain. Therefore, in order to produce useful machine learning models for binary analysis, it is necessary at a minimum to test, and likely necessary to train, these models on Windows data. We hope that will Assemblage fill the dataset gap to enable the application of modern efforts to Windows PE executables.
Comparing Assemblage and Preexisting Corpuses
Table 1 outlines a collection of binary corpuses from the literature, along with the largest corpus generated by Assemblage on top. The data set name is shown on the left, followed by the number of binaries, the number of functions (across all binaries), and the toolchains. The next column illustrates the dataset format: Assemblage provides a detailed source-to-binary mapping as part of the compilation process, and generates Program DataBase (PDB) files for use with reverse engineering tools such as IDA Pro. EMBER is a popular data set of Windows PE binaries, but the raw binaries are not public due to licensing restrictions; instead, EMBER provides only extracted features (Anderson and Roth, 2018). PassTell’s data set is also in the form of features, though its source is available and the basis of one of our case studies. While Windows PE binaries from XDA (generated by GHCC) are available, the total number of functions is relatively small (roughly 10k) (Pei et al., 2021); in more recent work using transformer learning, the XDA authors use Linux binaries gathered from open-source repositories to attain a sufficient amount of training data (Chen et al., 2022).
In sum, the vast bulk of work in modern binary analysis uses corpuses which are either (a) far too small for modern applications, or (b) consist exclusively of ELF binaries. Assemblage addresses the limitations of these existing datasets by providing large corpuses of both Linux and Windows binaries, complete with provenance and attribution to ensure compliance with licensing concerns and exposing the source-to-binary mapping.
3 Assemblage’s Design and Implementation
Figure 1: Assemblage architecture. A user downloads a corpus built via a worker pool on AWS, driven by the coordinator, which is configured via an interface accessed via secure shell.
We designed Assemblage as a cloud-based distributed system tailored to produce robust, reproducible, and high-quality binary corpuses. Figure 1 illustrates Assemblage’s principal components.
Distributed Architecture
Assemblage is designed to run as a distributed system on any cloud infrastructure, implementing a star topology system architecture. It consists of a central coordinator node, an adaptive worker pool, and a database interface. The system is designed to be resistant to component failure and endure high throughput. Assemblage provides rich APIs to enable wide flexibility in configuration, deployment, and extension.
Assemblage is driven by the coordinator node, which is responsible for configuration, task management, and database updating. Administrators communicate with Assemblage through a command-line interface provided by the coordinator. Then, a RabbitMQ message broker performs task management via a single-producer multi-consumer message queue. Finally, the coordinator dispatches tasks to specific channels of workers in the worker pool and, once a worker marks a task as complete, updates its database accordingly.
Assemblage’s worker nodes form an adaptive pool where each worker listens to one designated task channel. New workers can join the pool by registering with the coordinator and listening to a specific channel. Conversely, the coordinator can mark a worker as offline if it doesn’t send back a heartbeat signal within an allotted time period.
The modular design of Assemblage ensures that the unavailability of certain services does not result in a system-wide failure. The worker pool protects the system from cascading failures, so that the crash of one worker does not lead to a system-wide collapse. The message queue helps system communication remain stable as the worker pool scales. In practice, we have deployed up to 100 workers concurrently without issue.
Task Lifecycle
We will first describe our GitHub crawler, though Assemblage abstracts around input streams for source code, and we also include input sources for local (on-disc) repositories, Arch Linux’s AUR, and Microsoft’s vcpkg. The crawler begins by sending queries to GitHub’s APIs to discover new repositories. The coordinator then assembles results into build tasks, which instruct a worker to clone the repository and build it in a specific configuration. Workers listening to the channels will clone and build the binaries indicated in each task, and store the collected binaries in a designated storage location (typically an AWS S3 Bucket). Build workers are implemented by instantiating an abstract interface; we implemented builders for both Windows PE and Linux ELF binaries. On Windows, Assemblage uses the Microsoft Visual Studio toolchain, incorporating a set of automation scripts to overcome Visual Studio’s reliance on GUI-based user interfaces. On Linux, we implement workers for both GCC and Clang, using Docker to facilitate flexible and isolated deployment.
Tracking Build Provenance
Most prior work that built datasets of executables relied on disassembly tools to extract function information from binaries. For example, Wang et al. (2022), Ahn et al. (2022), and Marcelli et al. (2022) use IDA Pro to retrieve function information, while Du et al. (2022) uses objdump. Though significant engineering effort has gone into improving such disassembly tools over the years, they still do not work flawlessly, especially on Windows binaries. For example, IDA Pro can only infer function boundaries, but not names, on Windows PE binaries. Moreover, there is no guarantee as to the correctness of the recovered function information, especially when certain sections of information are stripped entirely(Wang et al., 2022). To address the shortcomings inherent in using tools such as IDA Pro, Assemblage is designed is carefully track build provenance, enabling the collection of ground truth compiler flags and settings.
Assemblage’s methods for recording build provenance are operating system dependent. For Linux binaries, we keep the compiled binary’s symbol table, unstripped, to preserve function information. This information is then added to the SQLite database corresponding to the dataset. For Windows binaries, in addition to collecting the binaries themselves, Assemblage also collects Program Database (PDB) files and other log files generated by the Microsoft Visual C++ compiler. These PDB files contain essential debug information that can be leveraged by binary analysis tools such as IDA Pro and Ghidra to associate source code with the bytes of compiled binaries and generate precise analysis outcomes. For the ease of generating training labels from PDB files, Assemblage processes PDB files utilizing Microsoft Debug Interface Access (DIA) SDK’s sample code “DIA2dump” (Microsoft, 2022a), then stores the files in exported Assemblage datasets. Additionally, the compilation metadata recorded in the PDB files is stored in the dataset’s SQLite database for portability.
By controlling the metadata passed to compiler during compilation, Assemblage ensures that the records in its SQLite table are accurate. To compare the utility of Assemblage’s metadata to that collected by previous methods, we passed a subset of Assemblage Windows binaries, containing a total of 23.7 million functions, to IDA Pro 8.2. IDA was unable to recover the names of nearly 10.7 million (45%) of those functions, instead returning a generic subroutine_location label (e.g. sub_10013040). The incidence of irretrievable function names in Assemblage’s metadata, by comparison, was $6\cdot 10^{-7}\%$. This demonstrates that Assemblage provides more complete and more accurate binary executable data than earlier dataset collection procedures.
Serializing “Recipes” for Reproducibility
The distribution of binary corpuses is complicated by the data size, licensing issues, and the need to accommodate application-specific features (e.g., control-flow graphs extracted via a commercial tool). To solve these issues, Assemblage supports distribution of datasets via “recipes,” compact Python bootstrap scripts that capture all information necessary to reconstruct a dataset. Given a recipe, users can replicate the state of a running Assemblage instance, as details about worker nodes and their associated instance types, build configurations, and repository URLs are all included. Recipes take advantage of Assemblage’s declarative API to enable wide flexibility in system operation. For example, via a small change in the recipe, Assemblage can be switched to either continuously crawl for new repositories, replicate an existing set of repositories, or build a set of repositories from a local disk. This allows sharing a smaller “recipe” over a larger data dump and a backup mechanism in the event of data corruption or hardware failure.
4 The Assemblage Datasets

### v2

@@ Abstract
Binary code is pervasive, and binary analysis is a key task in reverse engineering, malware classification, and vulnerability discovery. Unfortunately, while there exist large corpora of malicious binaries, obtaining high-quality corpora of benign binaries for modern systems has proven challenging (e.g., due to licensing issues). Consequently, machine learning based pipelines for binary analysis utilize either costly commercial corpora (e.g., VirusTotal) or open-source binaries (e.g., coreutils) available in limited quantities. To address these issues, we present Assemblage: an extensible distributed system that crawls, configures, and builds Windows PE binaries to obtain high-quality binary corpora suitable for training state-of-the-art models in binary analysis. We have run Assemblage on AWS over the past year, producing 890k Windows PE and 428k Linux ELF binaries across 29 configurations. Assemblage is designed to be both reproducible and extensible, enabling users to publish “recipes” for their datasets, and facilitating the extraction of a wide array of features. We evaluated Assemblage by using its data to train modern learning-based pipelines for compiler provenance and binary function similarity. Our results illustrate the practical need for robust corpora of high-quality Windows PE binaries in training modern learning-based binary analyses. Assemblage code is open sourced under the MIT license, and the dataset can be downloaded from https://assemblage-dataset.net/.
@@ 1 Introduction
Binary analysis, the task of understanding and analyzing an executable binary program, is extraordinarily labor-intensive. Practitioners often spend hours to weeks manually analyzing a single program [57]. The development of automated binary analysis tools has also proven to be extremely labor-intensive. Historically, such tools have relied on careful and judicious software engineering efforts to handle all corner cases, or used SAT solvers and similar data structures to resolve as many conditions as possible [45]. Given this high human effort, it is unsurprising that an increasing number of binary analysis methods use machine learning to help automate such efforts. This includes low-level tasks such as function disassembly [70], and high-level tasks like malware detection [59] or variable name identification [39, 14] which leverage these lower-level results.
@@
However, significant issues in data availability have hampered the development of machine learning-based binary analysis for decades [1, 56]. Disassembly and decompilation require a mapping between the original true source or assembly code and the corresponding bytes of binary instructions, but it is non-trivial to collect and maintain such a mapping. Malware detection requires access to a collection of “benign” programs to train on, but can not be legally distributed due to copyright and license concerns. The situation is particularly bad for Windows Portable Executable (PE) binaries: while Windows is the most frequent target of malware, modern binary analysis systems train almost exclusively on Linux ELF binaries. Toolchain identification requires knowledge of the entire compilation process and of a variety of compiler settings that are simply not recorded in the original file. Binary function similarity requires having a ground truth of functions that are semantically equivalent but written or compiled differently; most existing models rely on coding competition data that does not reflect the diversity of real-world code. The ad-hoc approaches taken by prior works to tackle these issues further complicate the reproducibility of these methods due to the infrastructure challenges in setting up an environment to build the same tools and perform complex data pre-processing [47, 54, 55, 25]. Moreover, authors often do not have the necessary licenses or do not make the code publicly available for others to extend the work.
@@
We introduce Assemblage, a tool that makes significant progress towards overcoming these challenges. Assemblage is a framework for crawling code hosting platforms, downloading open-source projects, compiling them with multiple different compiler versions and settings, and recording the necessary information to reconstruct byte-level mappings from the output back to the source code. By using open-source code targets from GitHub, we obtain greater diversity in programs, resolve licensing/distribution issues, and provide detailed ground truth to support all the aforementioned research areas. This alleviates the concerns with non-semantic preserving modifications [22] and the exorbitantly expensive cost and more limited diversity of binary re-writing [68, 11]. Assemblage is designed to be extensible so that additional sources of code, compilers, and feature extractors can be added for further functionality. Assemblage datasets can be further distributed as “recipes,” a list of repositories and settings so that issues with mixed-license repositories (e.g., GPL and BSD) do not make a binary-based distribution unlicensable. Finally, we have designed and tested the system to ensure the reproducibility of datasets built by Assemblage.
@@
(1) Assemblage, a cloud-based distributed system carefully engineered to automatically construct diverse corpora of high-quality Windows PE binaries while ensuring compliance with licensing and terms of service.
(2) The Assemblage dataset, consisting of Windows and Linux binaries with a total of 29 different combinations of CPU architecture, optimization level, compiler, compiler version, and other compiler flags.
(3) Three case studies that demonstrate Assemblage’s application to (a) compiler provenance detection, (b) function similarity identification and (c) function boundary identification. Our case studies demonstrate the need for Assemblage’s rich features, particularly for corpuses of Windows PE binaries.
2 Related Work
Dataset Binaries (#) Available data format Functions (#, K) Projects (#) OS Toolchain Supporting task L W G C M MD CP FS FN DC SPEC CPU2006 [5, 49] 981 PTY U 7 Ubuntu Dataset[6] 87,853 Binaries 88,000 22,040 PF[52] 2,132 N/A 1,362 5 PSI[52] 188, 253 N/A U 14,000 U U Nero[18] 542 Binaries U U U U U DIRE[39] 164,632 Binaries 3,196 U U U BinKit[36] 243,128 Binaries 75,231 51 Passtell[21] 552 Binaries 150 15 DIRT[13] 75,656 Binaries 998 U BinaryCorp -26M[66] 48,130 Binaries 25,877 9,819 BinBench[16] 1,127,479 Binaries 4,408 U LLM4- Decompile[65] U Binaries U 164 Assemblage 1,536,171 Binaries +PDB 783,694 220,792
Table 1: Comparison between Assemblage and related datasets. Prior datasets are often restricted in availability, do not have source-binary mappings, and are limited in platforms/toolchains, making it difficult to train and test for the diversity of real-world applications for binary analysis. Column names are abbreviated as; N/A: not available. L: Linux. W: Windows. G: GCC. C: Clang. M: MSVC. MD: malware detection. CP: compiler provenance. FS: function similarity. FN: function name/type recovery. DC: decompilation. PTY: The dataset is proprietary and for commercial license use. U: undisclosed or unknown.
@@ 2 Related Work
Some prior works have looked at crawling open-source repositories to build datasets  [43, 23, 10, 61] and leverage various build systems (such as DIRTY, XDA, and DIRE) [38]. Unfortunately, all these prior works focus on Linux binaries and build systems due to their ease of use. We found that nearly half of the transformer-based models rely on executables from standard Linux packages such as coreutils, binutils, diffutils, findutils, busybox, libcurl and openssl. With only one recent transformer-based model even testing on Windows PE Wu et al. [69]. The heavy Linux reliance [28, 31, 15, 62] and use of ubiquitous libraries [47, 36, 3, 17, 30] is problematic since most malware is Windows PE, and these utilities do not reflect the breadth of code types in the wild. Work that does focus on Windows data is almost exclusively distributed in a pre-processed and featurized form, forgoing the original binaries due to licensing issues [27, 4, 60, 73, 72, 53, 41], or purely malware and lacks any source-binary mappings [58, 34].
The lack of datasets with available source-to-binary mapping is also a major impediment to current research trends in the binary analysis space. Lessons/models from Natural Language Processing in developing BERT [19] and GPT [50] have shown that both architecture and data scaling are key factors to improved results [12, 67]. While many works are building base binary models that can be fine-tuned to perform a wide variety of tasks [41, 7, 37, 74], they predominantly use datasets that are too small or not diverse, as shown in Table 1 with a median of 3.7k binaries per LLM based model. Wang et al. [66] and Zhu et al. [74] are trained with the BinaryCorp dataset, which, at 48,130 (Linux) binaries, the largest prior work that still lacks Windows binaries, is not sufficiently diverse in binary behavior (as our results will show), and short of the 890k Windows PE and 428k Linux ELF binaries Assemblage provides. This is also true of works that tackle specific downstream tasks, including binary code similarity detection [66, 2, 24, 44], compiler provenance identification [29], vulnerability detection [71], and malware detection [40, 26, 69].
@@
With the increasing use of binary analysis in security-relevant settings, the absence of publicly available corpora of benign Windows binaries is a critical concern. The difficulty of obtaining Windows PE data in academic literature is prevalent in related domains like malware detection that rely on small datasets [34], effect sizes [51], and label noise [32, 33, 35], which generally depend on VirusShare and VirusTotal for data. An examination of files uploaded to VirusShare in 2023 indicates that 40.38% of the roughly 1.8 million malware samples are Windows binaries, whereas less than one percent are ELF binaries [58]. [^1] The other major file types are PDF, which accounts for roughly 25% of files, and Unknown, which accounts for 32% of files and captures files whose types could not be parsed. Files in the unknown category include data files, some scripts, ASC files, and source code. Given the prevalence of Windows executables, and the fact that Windows accounts for roughly seventy percent of global operating system market share [64], it is troubling that nearly all the recent literature on machine learning approaches to binary analysis uses Linux binaries as training data. This is especially worrying because, as we demonstrate in Section 4, models trained on Linux binaries do not necessarily generalize to the Windows domain. Therefore, in order to produce useful machine learning models for binary analysis, it is necessary at a minimum to test, and likely necessary to train, these models on Windows data.
3 The Assemblage Dataset
3.1 Dataset Generation Pipeline
The overview of Assemblage is illustrated in Figure 2 and Figure 2. The grey area indicated in Figure 2 shows Assemblage’s workflow, and it is implemented as a distributed system running on multiple servers for scalability. We also release the codes along with the dataset, and provides development and deployment documentation for it, more details are enclosed in our supplemental materials. First, the crawler begins by either sending queries to GitHub API to discover new repositories or crawl existing webpage hosting projects to obtain source codes. Then, source codes are cloned and modified to assign compiler flags such as optimization level and CPU architecture. On Windows, Assemblage uses the Microsoft Visual Studio tool chain, incorporating a set of automation scripts to overcome Visual Studio’s reliance on GUI-based user interfaces; while on Linux, workers are implemented for both GCC and Clang, using Docker to facilitate flexible and isolated deployment. After compiling is completed, all files will be collected to extract information (e.g., function address, source codes, comments) and stored in files. The last step is to compress all files for storage.
Figure 1: Assemblage’s build pipeline. Grey boxes are run in a parallel, distributed manner.
Figure 2: Assemblage architecture.
3.2 Dataset Details and Distribution